# neurips_2025.bib

### v1

@@ A.1 Real-world Experiments
Our CronusVLA adopts a multi-frame modeling strategy to improve the performance of VLA models, particularly in long-horizon tasks. We find that incorporating temporal information from multiple frames considerably enhances the model’s ability to perform multi-step execution, which is crucial for accurately completing complex, sequential instructions. As illustrated in Figure 7, we present a representative case from a real-world long-horizon task: pressing red, yellow, and green buttons in sequence. This task requires both instruction following and proper task decomposition. OpenVLA, as a single-frame model, struggles with these aspects and fails to execute the correct order, as shown in Figure 7 (A, Error 1). Furthermore, due to the presence of ambiguous states—where similar observations occur before and after button presses—OpenVLA cannot reliably distinguish between them, leading to repeated presses of the same button (A, Error 2). In contrast, CronusVLA demonstrates strong task decomposition and instruction-following capabilities, and its multi-frame temporal modeling provides critical context for resolving such ambiguities, enabling more robust and accurate execution in long-horizon scenarios.
@@
A.2 Simulation Experiments
Our model achieves a considerable performance breakthrough on the Put in Drawer task of SimplerEnv benchmark, reaching 64.8% and 65.1% in Visual Matching and Variant Aggregation, compared to baseline methods [openvla, spatialvla, robovlms, Tracevla], which achieve only 0%–30% success rates. As shown in Figure 8, we present a case study in the SimplerEnv setting with a Google Robot. In (A) and (B), OpenVLA struggles to fully open the drawer, blocking progress at the first step. In contrast, CronusVLA consistently completes all steps of the long-horizon task. In (C) and (D), although RoboVLMs can successfully open the top drawer initially, they fail to maintain the new execution goal, resulting in inaccurate actions and repeated collisions. This may stem from inadequate modeling capacity to adapt execution goals throughout long-horizon tasks. CronusVLA, by contrast, executes all steps smoothly and reliably.
Appendix B Additional Experiments
B.1 More Explorations on Multi-frame Training
@@
Single-frame pretraining considerably accelerates the convergence of multi-frame training. We compare the effectiveness of applying multi-frame post-training to a pretrained single-frame VLA model with training a multi-frame VLA entirely from scratch. To explore the differences under equal GPU budgets, we adopt identical settings and train a plain VLM model (without discrete embodied pretraining) for the same number of steps (50k for multi-frame modeling), as shown in Figure 9. In this comparison, Scratch (oxe) refers to training the plain VLM from scratch with multi-frame modeling, on the same OXE datasets as OpenVLA. Scratch (subset) denotes training the VLM with multi-frame modeling on the Fractal and Bridge-v2 datasets. The average SimplerEnv score of our method after training for 50k steps is 70.9, while models trained from scratch fail to converge within the limited training budget (10.4 v.s. 70.9 and 18.1 v.s. 70.9). We attribute this to the absence of prior embodied knowledge for training from scratch. Forcing the plain VLM to simultaneously learn embodied perception and multi-frame representations is an inherently challenging process within a limited training budget. Notably, even single-frame models, such as OpenVLA [openvla], typically require 28 to 40 epochs to pretrain on OXE, whereas our setting with 50k steps amounts to fewer than 5 epochs, making training from scratch particularly challenging. Our method can leverage pretrained on-the-shelf VLA checkpoints (e.g., [openvla, spatialvla, belkhale2024minivla]) in the community, making it a more practical and resource-efficient alternative.
@@
Pretraining with multi-frame regularization. We also investigate the impact of our multi-frame regularization on training a multi-frame VLA entirely from scratch. Without this regularization, as shown by the SimplerEnv scores of Scratch (subset) w/o reg. and Scratch (oxe) w/o reg. in Figure 9, both settings exhibit lower average success rates and a tendency toward slower convergence compared to their regularized counterparts. These results further highlight the effectiveness of multi-frame regularization in facilitating stable and efficient multi-frame training.
@@
Figure 9: Convergence trends of training multi-frame VLA entirely from scratch, based on CronusVLA 7B (6 past frames) and tested on SimplerEnv.
Table 6: Ablation on motion feature types.
@@
final 61.4 56.6 40.6 52.9
B.2 Ablation on Motion Feature Types
Building on the CronusVLA 0.5B model with a minimal frame count (1 past frame, a total of 2), we investigate the impact of different motion feature representations on performance, focusing on settings with limited temporal context. Specifically, we compare three configurations: (1) final, which uses the last-layer of a single learnable feature placed after all visual and language tokens; (2) visual + final, which uses two learnable features, and one inserted after the final visual token and the other after all tokens, all features are extracted from the last layer; and (3) multi-layer final, which extracts the same position’s features from three different LLM layers (layers 24, 16, and 8 out of 24). As shown in Table 6, multi-layer final improves performance but increases model complexity, reducing overall usability. Additionally, incorporating extra visual tokens degrades performance, possibly due to misalignment with the VLM’s pretrained causal structure, increasing optimization difficulty. Balancing performance and efficiency, we adopt the final strategy in our main model.
B.3 Ablation on Post-training Data
As shown in Table 7, we evaluate the performance of our model in SimplerEnv under different post-training datasets. All experiments are conducted using the CronusVLA 0.5B model with a total two-frame input. We consider three configurations: (1) training on Bridge-v2 [bridgev2] and evaluating on the WidowX robot, (2) training on Fractal [RT1] and evaluating on the Google robot, and (3) joint training on both datasets with evaluation on both platforms. Results indicate that joint training consistently improves performance across robots. We attribute this to the complementary nature of the datasets, which facilitates learning more robust and generalized representations, thereby mitigating overfitting associated with single-dataset training.
Table 7: Ablation on post-training data.
Fractal Bridge Google VM Google VA WidowX VM
@@ B.3 Ablation on Post-training Data
61.4 56.6 40.6
B.4 Hyperparameter Sensitivity Analysis of Action Adaptation
We investigate the impact of two key hyperparameters in our action adaptation module: the number of top-k nearest neighbors ($K$) and the similarity threshold ($B$), which together govern how reference actions are aggregated and filtered based on quality. In Figure 10 (1) and (2), we show how varying top $K$ affects performance on LIBERO-Long and the LIBERO average across all tasks. In Figure 10 (3) and (4), we illustrate the effect of different threshold $B$ values. The results (1) and (2) suggest that too few sampled actions may lack robustness, while too many can reduce informativeness. A properly chosen top $K$ enables the model to collect meaningful guidance. Similarly, a low threshold $B$ admits low-quality sampled actions that degrade performance, while a high threshold $B$ may overly restrict the pool, limiting access to useful references.
Figure 10: Hyperparameter sensitivity analysis of action adaptation. We study the impact of two key hyperparameters in the action adaptation module across LIBERO tasks, including top-k nearest neighbors and the similarity threshold.
B.5 Detailed Main Result in SimplerEnv
We report detailed evaluation results for all SimplerEnv settings, focusing on key baselines for comparison. Table 8 presents results on the Google Robot across both Visual Matching and Variant Aggregation, covering four tasks. Extended evaluations include coke can manipulation (horizontal, vertical, and upright grasping) and drawer manipulation (opening and closing). Visualization examples are provided in Figure 18 and Figure 19. Additionally, Table 9 reports results on the WidowX Robot, comparing our models against representative baselines. More visualization results are shown in Figure 20. Metrics include grasp success rate and overall task success. Averaged across all tasks, both CronusVLA 0.5B and CronusVLA 7B outperform prior baselines in most cases, demonstrating the effectiveness of our training strategy and model architecture in achieving robust performance and generalization.
@@ B.5 Detailed Main Result in SimplerEnv
Horizontal Laying Vertical Laying Standing Average Average Open Close Average Average Average
Visual Matching RT-1-X 82.0 33.0 55.0 56.7 31.7 29.6 89.1 59.7 21.3 42.4
@@
CronusVLA (0.5B) 90.0 100.0 98.0 96.0 93.0 46.3 54.6 50.5 42.6 70.5
CronusVLA (7B) 95.0 94.0 98.0 95.7 76.0 62.0 93.5 77.8 64.8 78.6
Variant Aggregation RT-1-X 56.9 20.4 69.8 49.0 32.3 6.9 51.9 29.4 10.1 30.2
@@
CronusVLA (0.5B) 54.2 45.8 54.2 33.3 58.3 0.0 79.2 79.2 39.6
CronusVLA (7B) 75.0 66.7 79.2 54.2 41.7 20.8 100.0 100.0 60.4
B.6 More Visualization Results in LIBERO Benchmark
We illustrate some visualization results of four task suites in Figure 21, including LIBERO Spatial, LIBERO Object, LIBERO Goal, and LIBERO Long.
Appendix C Training Details
@@
The three training stages, including single-frame pretraining, multi-frame post-training, and finetuning, are illustrated in the  Figure 11. We focus on the multi-frame post-training and finetuning in this paper.
C.1 Pretraining Details
@@
For the 7B model, we follow the training setup and protocol of OpenVLA [openvla] by leveraging 27 datasets from Open X-Embodiment [oxe]. The same procedure is applied to the 0.5B model based on the VLM of [belkhale2024minivla]. Training is performed by minimizing the cross-entropy loss between the predicted token distribution and ground truth tokens. For more details, please refer to [openvla].
C.2 Post-training Details
@@
Our models are post-trained on cross-embodiment datasets, including Fractal[RT2] and Bridge-v2[bridgev2]. We follow the data processing pipeline of Octo [octo] and OpenVLA [openvla], including RLDS preprocessing and augmentations such as random resized cropping, and adjustments to brightness, contrast, saturation, and hue. To enhance generalization, the data loader randomly samples across datasets, trajectories, and time steps. All components, including the vision encoder, projection layer, language model, and decoder, are trained end-to-end by minimizing the mean squared error between predicted and ground-truth diffusion noise. Training employs a 100-step diffusion noise schedule. The language model does not generate text tokens autoregressively; we omit text prediction cross-entropy loss and discard the language head.
@@
CronusVLA 7B is initialized from the pretrained single-frame 7B VLA model. Post-training is conducted on 64 A100 GPUs for approximately 50k gradient steps, with a total batch size of 512 and default 6-frame past observation sequences. We use AdamW with a learning rate of 4e-5 and a linear scheduler, without weight decay or warm-up. PyTorch FSDP is applied to reduce memory usage. Due to the inherent instability of diffusion-based training, which is sensitive to randomness, we select the best-performing checkpoint between 45k and 55k steps, evaluated every 2.5k steps. CronusVLA 0.5B utilizes Qwen2.5 0.5B [qwen2] as the LLM backbone, SigLIP [siglip] and DINOv2 [dinov2] as the vision encoder. Post-training is performed on 32 A100 GPUs, about 40k gradient steps, with a batch size of 1024 and default 3-frame past observation sequences. We also use AdamW with a learning rate of 4e-5 and a linear scheduler, without weight decay or warm-up.
C.3 Finetuning Details
@@
Building upon the post-trained models, we further finetune on expert demonstrations using an action adaptation strategy. As shown in Figure 12, the process consists of two phases. In the first phase, the post-trained model is adapted to the specific task setup (such as the task suites of LIBERO) by finetuning the whole model (except for Adapter). CronusVLA-7B is finetuned for 10K steps across all tasks. In the second phase, the adapted model is used to construct a retrieval matrix from expert demonstrations, which guides further training. During this phase, the Vision-Language Backbone and Modulator are frozen, and only the Action Adapter and Decoder are trained for an additional 5K–12K steps. Training is conducted with a batch size of 256 on 16 A100 GPUs and a total 15K–22K steps, using a learning rate of 2e-5 and a fixed sequence length of 3-frame past observations.
Appendix D Model Details
Figure 12: Detailed Design. (a) illustrates the two-stage finetuning process. In the first stage, the Vision-Language Backbone, Modulator, and Decoder are trained jointly to adapt to the task. In the second stage, the Vision-Language Backbone and Modulator are frozen to stabilize the learned representations, and only the Adapter and Decoder are fine-tuned. (b) presents the feature chunking mechanism employed during inference, where the model processes a sequence consisting of three previous frames and one current frame. (c) demonstrates the action adaptation mechanism, in which the reference action is injected into the decoder.
D.1 Model Architecture
@@
As shown in Table 10, we detail the model architecture. The language model is either LLaMA2 or Qwen2.5, and the vision encoder combines SigLIP and DINOv2, taking 224×224 RGB images as input. A linear layer serves as the cross-modal projector (MM projector). These components comprise the core VLA module, which processes $B\cdot F$ individual samples. The decoder integrates multiple inputs: ground-truth actions, timestep encodings, and multi-frame motion features from the VLA. Actions and timesteps are embedded into a 768-dimensional space via the Action and Timestep Embedders, respectively. The Action Adapter is activated only during finetuning. Motion features are modulated by the Modulator, and the decoder consists of 12 layers of Self-Attention and Cross-Attention, facilitating interaction between noised actions and motion features.
D.2 Feature Chunking
The core logic of feature chunking is depicted in Figure 12 (b). At inference, feature chunking accelerates prediction by caching previously computed motion features. To maintain a consistent feature length M at each step, a first-in-first-out queue is employed. For early timesteps (t < M) with insufficient historical features, we adopt a standard low-level policy strategy by padding the buffer with repeated first-frame features. As similar padding is introduced during training via stochastic sampling, this method does not degrade performance.
D.3 Action Adapter
To guide final action generation, reference actions, which are retrieved based on similarity in the modulated feature space, are incorporated into the decoder as lightweight plugins, as shown in  Figure 12 (c). These reference actions act as auxiliary cues rather than strict supervision, given the challenge of obtaining precise ground-truth actions via retrieval. Empirically, we adopt a straightforward approach: the reference action is concatenated with the noised action and passed through an MLP to project it into the decoder’s latent space. The resulting embedding is added to the input noised actions as an auxiliary signal, applied only during the initial processing and not propagated across subsequent decoder layers. For each finetuning demonstration dataset, specifically the four LIBERO tasks, a retrieval matrix is constructed. Following the random sampling strategy used during training, we restrict the retrieval similarity threshold to 0.6. To utilize the ground-truth actions during training, the most similar action is explicitly masked. The retrieval process is summarized in  Algorithm 1.
Table 10: Model architecture components with model size, layer numbers, and tensor shapes. Notation: $B$ = batch size, $F$ = total number of past and current frames, $T$ = text token count, $A$ = action sequence length.
@@ D.3 Action Adapter
Input: $\mathcal{F}_{t}$: Modulated features at time $t$
Input: $\mathcal{D}=\{(\mathbf{A}_{i},\mathbf{F}_{i})\}_{i=1}^{N}$: Dataset of expert actions and modulated features
Input: $K$: Top-$K$ nearest neighbors
Input: $T$: Softmax temperature
Input: $B$: Similarity threshold
Output: Retrieved action $\hat{\mathbf{A}}_{t}$
Initialize: memory $\mathcal{M}\leftarrow\emptyset$ ;
AddToMemory($(\{\mathbf{A}_{i}\},\{\mathbf{F}_{i}\})$:
foreach $i$ do
       $\mathcal{M}\leftarrow\mathcal{M}\cup\{(\mathbf{A}_{i},\mathbf{F}_{i})\}$ ;
        // Append pair to memory
DeduplicateMemory():
Initialize hash set $\mathcal{H}\leftarrow\emptyset$, deduplicated memory $\mathcal{M}^{\prime}\leftarrow\emptyset$ ;
foreach $(\mathbf{A},\mathbf{F})\in\mathcal{M}$ do
       $h\leftarrow\texttt{Hash}(\mathbf{A},\mathbf{F})$ ;
       if $h\notin\mathcal{H}$ then
             $\mathcal{M}^{\prime}\leftarrow\mathcal{M}^{\prime}\cup\{(\mathbf{A},\mathbf{F})\}$ ;
             $\mathcal{H}\leftarrow\mathcal{H}\cup\{h\}$ ;
$\mathcal{M}\leftarrow\mathcal{M}^{\prime}$ ;
Save/Load:
Save$(\mathcal{M},\text{path})$ ;
Load$(\text{path})\rightarrow\mathcal{M}$ ;
Upon loading: call DeduplicateMemory(), BuildIndex() ;
BuildIndex():
$\mathbf{A}_{\text{all}}\leftarrow\texttt{Stack}(\mathbf{A}\in\mathcal{M})$ ;
$\mathbf{F}_{\text{all}}\leftarrow\texttt{Normalize}(\texttt{Stack}(\mathbf{F}\in\mathcal{M}))$ ;
return $\mathbf{A}_{\text{all}},\mathbf{F}_{\text{all}}$ ;
Retrieve$(\mathcal{F}_{t})$:
$\mathcal{F}_{t}\leftarrow\texttt{Normalize}(\mathcal{F}_{t})$ ;
$\mathbf{S}\leftarrow\mathcal{F}_{t}\cdot\mathbf{F}_{\text{all}}^{\top}$ ;
  // Similarity matrix

### v2

@@ B.2 Real-world Experiments
Our CronusVLA adopts a multi-frame modeling strategy to improve the performance of VLA models, particularly in long-horizon tasks. We find that incorporating temporal information from multiple frames considerably enhances the model’s ability to perform multi-step execution, which is crucial for accurately completing complex, sequential instructions. As illustrated in Fig.12, we present a representative case from a real-world long-horizon task: pressing red, yellow, and green buttons in sequence. This task requires both instruction following and proper task decomposition. OpenVLA, as a single-frame model, struggles with these aspects and fails to execute the correct order, as shown in Fig.12 (A, Error 1). Furthermore, due to the presence of ambiguous states, where similar observations occur before and after button presses, OpenVLA cannot reliably distinguish between them, leading to repeated presses of the same button (A, Error 2). In contrast, CronusVLA demonstrates strong task decomposition and instruction-following capabilities, and its multi-frame temporal modeling provides critical context for resolving such ambiguities, enabling more robust and accurate execution in long-horizon scenarios.
@@
Appendix C Additional Experiments
C.1 More Explorations on Multi-frame Training
@@
Single-frame pretraining considerably accelerates the convergence of multi-frame training. We compare the effectiveness of applying multi-frame post-training to a pretrained single-frame VLA model with training a multi-frame VLA entirely from scratch. To explore the differences under equal GPU budgets, we adopt identical settings and train a plain VLM model (without discrete embodied pretraining) for the same number of steps (50k for multi-frame modeling), as shown in Fig.13. In this comparison, Scratch (oxe) refers to training the plain VLM from scratch with multi-frame modeling, on the same OXE datasets as OpenVLA. Scratch (subset) denotes training the VLM with multi-frame modeling on the Fractal and Bridge-v2 datasets. The average SimplerEnv score of our method after training for 50k steps is 70.9, while models trained from scratch fail to converge within the limited training budget (10.4 v.s. 70.9 and 18.1 v.s. 70.9). We attribute this to the absence of prior embodied knowledge for training from scratch. Forcing the plain VLM to simultaneously learn embodied perception and multi-frame representations is an inherently challenging process within a limited training budget. Notably, even single-frame models, such as OpenVLA (Kim et al. 2024), typically require 28 to 40 epochs to pretrain on OXE, whereas our setting with 50k steps amounts to fewer than 5 epochs, making training from scratch particularly challenging. Our method can leverage pretrained off-the-shelf VLA checkpoints (e.g., (Kim et al. 2024; Qu et al. 2025; Belkhale and Sadigh 2024)) in the community, making it a more practical and resource-efficient alternative.
@@
Pretraining without multi-frame regularization. We also investigate the impact of our multi-frame regularization on training a multi-frame VLA entirely from scratch. Without this regularization, as shown by the SimplerEnv scores of Scratch (subset) w/o reg. and Scratch (oxe) w/o reg. in Fig.13, both settings exhibit lower average success rates and a tendency toward slower convergence compared to their regularized counterparts. These results further highlight the effectiveness of multi-frame regularization in facilitating stable and efficient multi-frame training.
@@
Figure 13: Single-frame pretraining and multi-frame regularization accelerate the convergence of multi-frame training. Convergence trends of training multi-frame VLA are shown, based on CronusVLA 7B (6 past frames) and SimplerEnv.
C.2 Ablation on Types of Learnable Feature
Building on the CronusVLA 0.5B model with a minimal frame count (1 past frame, a total of 2), we investigate the impact of different learnable feature representations on performance, focusing on settings with limited temporal context. Specifically, we compare three configurations: (1) final, which uses the last-layer of a single learnable feature placed after all visual and language tokens; (2) visual + final, which uses two learnable features, and one inserted after the final visual token and the other after all tokens, all features are extracted from the last layer; and (3) multi-layer final, which extracts the same position’s features from three different LLM layers (layers 24, 16, and 8 out of 24). As shown in Tab.12, multi-layer final improves performance but increases model complexity, reducing overall usability. Additionally, incorporating extra visual tokens degrades performance, possibly due to misalignment with the VLM’s pretrained causal structure, increasing optimization difficulty. Balancing performance and efficiency, we adopt the final strategy in our main model.
@@ C.2 Ablation on Types of Learnable Feature
[…]
Table 12: Ablation on learnable feature types.
C.3 Ablation on Post-training Data
As shown in Tab.13, we evaluate the performance of our model in SimplerEnv under different post-training datasets. All experiments are conducted using the CronusVLA 0.5B model with a total two-frame input. We consider three configurations: (1) training on Bridge-v2 (Walke et al. 2023) and evaluating on the WidowX robot, (2) training on Fractal (Brohan et al. 2022) and evaluating on the Google robot, and (3) joint training on both datasets with evaluation on both platforms. Results indicate that joint training consistently improves performance across robots. We attribute this to the complementary nature of the datasets, which facilitates learning more robust and generalized representations, thereby mitigating overfitting associated with single-dataset training.
Fractal
Bridge
Google VM Google VA WidowX VM
@@ C.3 Ablation on Post-training Data
[…]
Table 13: Ablation on post-training data.
C.4 Detailed Main Result in SimplerEnv
We report detailed evaluation results for all SimplerEnv settings, focusing on key baselines for comparison. Tab.14 presents results on the Google Robot across both Visual Matching and Variant Aggregation, covering four tasks. Extended evaluations include coke can manipulation (horizontal, vertical, and upright grasping) and drawer manipulation (opening and closing). Visualization examples are provided in Fig.21 and Fig.22. Additionally, Tab.15 reports results on the WidowX Robot, comparing our models against representative baselines. More visualization results are shown in Fig.23. Metrics include grasp success rate and overall task success. Averaged across all tasks, both CronusVLA 0.5B and CronusVLA 7B outperform prior baselines in most cases, demonstrating the effectiveness of our training strategy and model architecture in achieving robust performance and generalization.
@@ C.4 Detailed Main Result in SimplerEnv
Horizontal
Laying
Vertical
Laying
Standing Average Average Open Close Average Average Average
Visual
Matching
RT-1-X 82.0 33.0 55.0 56.7 31.7 29.6 89.1 59.7 21.3 42.4
@@
[…]
Variant
Aggregation
RT-1-X 56.9 20.4 69.8 49.0 32.3 6.9 51.9 29.4 10.1 30.2
@@
[…]
C.5 More Results in LIBERO Benchmark
Methods Spatial Object Goal Long (10) Ave.
Type I: without state, without wrist view
Diffusion Policy (Chi et al. 2023) 78.5 $\pm$ 1.1 87.5 $\pm$ 0.7 73.5 $\pm$ 1.2 64.8 $\pm$ 1.3 76.1 $\pm$ 0.7
OpenVLA (Kim et al. 2024) 84.7 $\pm$ 0.9 88.4 $\pm$ 0.8 79.2 $\pm$ 1.0 53.7 $\pm$ 1.3 76.5 $\pm$ 0.6
TraceVLA (Zheng et al. 2024) 84.6 $\pm$ 0.2 85.2 $\pm$ 0.4 75.1 $\pm$ 0.3 54.1 $\pm$ 1.0 74.8 $\pm$ 0.5
SpatialVLA (Qu et al. 2025) 88.2 $\pm$ 0.5 89.9 $\pm$ 0.7 78.6 $\pm$ 0.6 55.5 $\pm$ 1.0 78.1 $\pm$ 0.7
Dita (Hou et al. 2025) 84.2 96.3 85.4 63.8 82.4
[…]
Type II: with state, with wrist view
GR00T-N1 (Bjorck et al. 2025) 94.4 $\pm$ 0.9 97.6 $\pm$ 1.0 93.0 $\pm$ 1.2 90.6 $\pm$ 1.0 93.9 $\pm$ 1.1
$\pi_{0}$ (Black et al. 2024) 96.8 $\pm$ 0.8 98.8 $\pm$ 0.9 95.8 $\pm$ 1.1 85.2 $\pm$ 1.2 94.2 $\pm$ 0.9
$\pi_{0}$-FAST (Pertsch et al. 2025) 96.4 $\pm$ 0.7 96.8 $\pm$ 0.7 88.6 $\pm$ 1.0 60.2 $\pm$ 1.4 85.5 $\pm$ 1.0
OpenVLA-OFT (Kim, Finn, and Liang 2025) 97.6 $\pm$ 0.9 98.4 $\pm$ 0.8 97.9 $\pm$ 1.0 94.5 $\pm$ 1.3 97.1 $\pm$ 0.6
$\pi_{0.5}+KI$ (from scratch) (Intelligence et al. 2025) 96.6 97.2 94.6 84.8 93.3
$\pi_{0.5}+KI$ (from generalist model) (Intelligence et al. 2025) 98.0 97.8 95.6 85.8 94.3
[…]
Table 16: Detailed LIBERO Benchmark Results. We report the success rate (SR) and standard error for each method across four task suites, averaged over three random seeds with 500 trials. To ensure fair comparison under different settings, we evaluate two categories of VLA models: Type I, which relies solely on the instruction and third-person view (e.g., OpenVLA, SpatialVLA); and Type II, which additionally incorporates state and wrist-view inputs (e.g., $\pi_{0}$, OpenVLA-OFT).
Detailed results on LIBERO are listed in Tab.16. Our model achieves performance comparable to the prior SOTA OpenVLA-OFT, despite not using state inputs or additional optimization, and requiring substantially fewer training steps (30K vs. 150K). We illustrate some visualization results of four task suites in Fig.24, including LIBERO Spatial, LIBERO Object, LIBERO Goal, and LIBERO Long.
Appendix D Training Details
@@
The training stages, including single-frame pretraining, multi-frame post-training, and finetuning, are illustrated in the Fig.14.
D.1 Pretraining Details
@@
For the 7B model, we follow the training setup and protocol of OpenVLA (Kim et al. 2024) by leveraging 27 datasets from Open X-Embodiment (O’Neill et al. 2024). The same procedure is applied to the 0.5B model based on the VLM of (Belkhale and Sadigh 2024). Training is performed by minimizing the cross-entropy loss between the predicted token distribution and ground truth tokens. For more details, please refer to (Kim et al. 2024).
D.2 Post-training Details
@@
Our models are post-trained on cross-embodiment datasets, including Fractal(Brohan et al. 2023) and Bridge-v2(Walke et al. 2023). We follow the data processing pipeline of Octo (Octo Model Team et al. 2024) and OpenVLA (Kim et al. 2024), including RLDS preprocessing and augmentations such as random resized cropping, and adjustments to brightness, contrast, saturation, and hue. To enhance generalization, the data loader randomly samples across datasets, trajectories, and time steps. All components, including the vision encoder, projection layer, language model, and decoder, are trained end-to-end by minimizing the mean squared error between predicted and ground-truth diffusion noise. Training employs a 100-step diffusion noise schedule. The language model does not generate text tokens autoregressively; we omit text prediction cross-entropy loss and discard the language head.
@@
CronusVLA 7B is initialized from the pretrained single-frame 7B VLA model. Post-training is conducted on 64 A100 GPUs for approximately 50k gradient steps, with a total batch size of 512 and default 6-frame past observation sequences. We use AdamW with a learning rate of 4e-5 and a linear scheduler, without weight decay or warm-up. PyTorch FSDP is applied to reduce memory usage. Due to the inherent instability of diffusion-based training, which is sensitive to randomness, we select the best-performing checkpoint between 45k and 55k steps, evaluated every 2.5k steps. CronusVLA 0.5B utilizes Qwen2.5 0.5B (Yang et al. 2024b) as the LLM backbone, SigLIP (Zhai et al. 2023) and DINOv2 (Oquab et al. 2023) as the vision encoder. Post-training is performed on 32 A100 GPUs, about 40k gradient steps, with a batch size of 1024 and default 3-frame past observation sequences. Action chunking (Zhao et al. 2023) is employed by following the settings of (Li et al. 2024a) with a chunking length of 16 and an adaptive ensemble strategy. We also use AdamW with a learning rate of 4e-5 and a linear scheduler, without weight decay or warm-up.
D.3 Finetuning Details
@@
Building upon the post-trained models, we further finetune on expert demonstrations using a full finetuning strategy. The post-trained weights are adapted to the specific task setup (such as each task suite of LIBERO). CronusVLA-7B is finetuned for 6K-30K steps. Training is conducted with a batch size of 256 on 16 A100 GPUs, using a constant learning rate of 2e-5 and a fixed sequence length of 3-frame past observations, and the action chunking length is 8.
Appendix E Model Details
E.1 Model Architecture
@@
As shown in Tab.17, we detail the model architecture. The language model is either LLaMA2 or Qwen2.5, and the vision encoder combines SigLIP and DINOv2, taking 224×224 RGB images as input. A linear layer serves as the cross-modal projector (MM projector). These components comprise the core VLA module, which processes $B\cdot F$ individual samples. The decoder integrates multiple inputs: ground-truth actions, timestep encodings, and multi-frame learnable features from the VLA. Actions and timesteps are embedded into a 768-dimensional space via the Action and Timestep Embedders, respectively. The Action Adapter is activated only during finetuning. Learnable features are modulated by the Modulator, and the decoder consists of 12 layers of Self-Attention and Cross-Attention, facilitating interaction between noised actions and learnable features. MLP in our model includes two linear layers with the GELU function. Modulator includes a DIV function(as shown in Eq.2) and a learnable positional embedding (as shown in Fig.2 (d)).
@@
Table 17: Model architecture components with model size, layer numbers, and tensor shapes. Notation: $B$ = batch size, $F$ = total number of past and current frames, $T$ = text token count, AA = action sequence length.
E.2 Feature Chunking
The core logic of feature chunking is depicted in Fig.15 (b). At inference, feature chunking accelerates prediction by caching previously computed learnable features. To maintain a consistent feature length M at each step, a first-in-first-out queue is employed. For early timesteps ($t<M$) with insufficient historical features, we adopt a standard low-level policy strategy by padding the buffer with repeated first-frame features. As similar padding is introduced during training via stochastic sampling, this method does not degrade performance.
Figure 15: Details of the feature chunking mechanism. During inference, the model processes a sequence consisting of several previous frames and one current frame.
E.3 Efficiency Analysis
Our model improves both performance and inference efficiency, and supports limited-cost training. An additional comparison is shown in Table 18. Our single-frame pretraining can be replaced by reusing or sharing public checkpoints (e.g., OpenVLA (Kim et al. 2024)), while the post-training stage requires only $\sim$50K steps, which is significantly fewer than other VLA counterparts. Our design supports not only a compact 0.5B model but also a larger 7B model. As the history length of CronusVLA-7B increases from 2 to 7, the inference speed remains constant (8.7 Hz) in Fig.5. The proposed multi-frame regularization employs no_grad for past frames with only one forward pass, ensuring efficient and scalable training.
Method
Frame
Number
Success